OpenStudy (kirbykirby):

Is this an appropriate way to show that 2x^2 - 4xy + 4y^2 >= 0?

OpenStudy (kirbykirby):

If x = y, then 2x^2 - 4y^2 + 4y^2 >= 0 2x^2 + 4y^2 >= 4y^2.. which is obvious and if y=x, 2x^2 - 4x^2 + 4y^2 >=0 2x^2 + 4y^2 >= 4x^2 4y^2 >= 2x^2

OpenStudy (anonymous):

Can I suggest another method?

OpenStudy (kirbykirby):

^Sure

OpenStudy (anonymous):

2x^2 - 4xy + 4y^2 x^2 + x^2 -2*x*2y +(2y)^2 x^2+(x-y)^2

OpenStudy (anonymous):

Now, for any values of x and y, x^2>=0 AND (x-y)^2>=0 Thus, x^2+(x-y)^2>=0

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

AND SORRY I DONT THINK WHAT U DID IS CORRECT

OpenStudy (kirbykirby):

Oh ok thanks :)

OpenStudy (anonymous):

WELCOME